for powerdot – slide 1

“Telling More“: a 2006 NASA SARP project

Tim Menzies

Lane Department of Computer Science and Electrical Engineering

West Virginia University

tim@menzies.us

May 29, 2006

Outline

Outline This section has a slide for powerdot – slide 2

Outline

This section has a slide Slide 2 Slide 3

outline

Outline This section has a slide for powerdot – slide 3

Outline

This section has a slide Slide 2 Slide 3

This section has a slide

Outline

Slide 2

This section has a slide Slide 2 Slide 3 for powerdot – slide 4 Outline

This section has a slide Slide 2 Slide 3 for powerdot – slide 5

Here is the binomium formula.

(a + b)^{n} = ∑_{k=0}^{n} \binom{n}{k}a^{n−k}b^{k} (1)

We will prove formula (1) on the blackboard. Click the title of this slide to switch to the black slide.

Outline

This section has a slide Slide 2 Slide 3 for powerdot – slide 5

Here is the binomium formula.

(a + b)^{n} = ∑_{k=0}^{n} \binom{n}{k}a^{n−k}b^{k} (1)

We will prove formula (1) on the blackboard. Click the title of this slide to switch to the black slide.

Outline

This section has a slide Slide 2 Slide 3 for powerdot – slide 5

Here is the binomium formula.

(a + b)^{n} = ∑_{k=0}^{n} \binom{n}{k}a^{n−k}b^{k} (1)

We will prove formula (1) on the blackboard. Click the title of this slide to switch to the black slide.

Outline

This section has a slide Slide 2 Slide 3 for powerdot – slide 6

This happens…

when you change…

the type of itemize.

Outline

This section has a slide Slide 2 Slide 3 for powerdot – slide 6

This happens…

when you change…

the type of itemize.

Outline

This section has a slide Slide 2 Slide 3 for powerdot – slide 6

This happens…

when you change…

the type of itemize.

Example 2  A hidden wide section  ^{for powerdot – slide 7}

Outline

This section has a slide

A hidden wide section Slide 4 Slide 5 one 5 for powerdot – slide 8

We only treat this material…

if we have some time left.

But don’t hesitate…

to read it yourself.

Outline

This section has a slide

A hidden wide section Slide 4 Slide 5 one 5 for powerdot – slide 8

We only treat this material…

if we have some time left.

But don’t hesitate…

to read it yourself.

Outline

This section has a slide

A hidden wide section Slide 4 Slide 5 one 5 for powerdot – slide 8

We only treat this material…

if we have some time left.

But don’t hesitate…

to read it yourself.

Outline

This section has a slide

A hidden wide section Slide 4 Slide 5 one 5 for powerdot – slide 8

We only treat this material…

if we have some time left.

But don’t hesitate…

to read it yourself.

Slide 5

Example 2 for powerdot – slide 9

This wide slide can contain more material as it is wider and does not have a table of contents.

Example 2  one  ^{for powerdot – slide 10}

Outline

This section has a slide

A hidden wide section Slide 4 Slide 5 one 5 for powerdot – slide 11

round 0: Cost Benefit  1  2  3  4 Totals 12 6  1  2  1  4 0 13  17  15  4  49 -6 10  12  4  1  27 -12  4  6  2  12 -18  3  2  1  1  7 Totals 30  38  24  6  98